### v1

# Compressed Feature Quality Assessment: Dataset and Baselines

Abstract.
The widespread deployment of large models in resource-constrained environments has underscored the need for efficient transmission of intermediate feature representations. In this context, feature coding, which compresses features into compact bitstreams, becomes a critical component for scenarios involving feature transmission, storage, and reuse. However, this compression process introduces inherent semantic degradation that is notoriously difficult to quantify with traditional metrics. To address this, this paper introduces the research problem of Compressed Feature Quality Assessment (CFQA), which seeks to evaluate the semantic fidelity of compressed features. To advance CFQA research, we propose the first benchmark dataset, comprising 300 original features and 12000 compressed features derived from three vision tasks and four feature codecs. Task-specific performance drops are provided as true semantic distortion for the evaluation of CFQA metrics. We assess the performance of three widely used metrics (MSE, cosine similarity, and Centered Kernel Alignment) in capturing semantic degradation. The results underscore the representativeness of the dataset and highlight the need for more refined metrics capable of addressing the nuances of semantic distortion in compressed features. To facilitate the ongoing development of CFQA research, we release the dataset and all accompanying source code at https://github.com/chansongoal/Compressed-Feature-Quality-Assessment. This contribution aims to advance the field and provide a foundational resource for the community to explore CFQA.
Compressed Feature Quality Assessment, Coding for Machines, Feature Coding
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Multimedia databasesccs: Computing methodologies Image compressionccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
Figure 1. Exemplar application scenarios of compressed feature quality assessment.
1. Introduction
The rapid deployment of large foundation models (e.g., DINOv2, LLaMA3) across distributed and resource-constrained environments has created a growing need to transmit intermediate features rather than raw signals. Feature coding, which compresses these intermediate representations, plays a vital role in enabling scalable, privacy-preserving, and efficient systems. Unlike traditional image or video coding that prioritizes perceptual quality, feature coding targets the preservation of task-relevant semantics embedded in deep representations.
However, feature coding inevitably introduces semantic degradation: a loss of discriminative information that can compromise downstream performance. This degradation is fundamentally different from pixel-level distortions and often cannot be captured by conventional distortion metrics such as MSE or PSNR. While task accuracy is a more reliable indicator of semantic quality, it is impractical for deployment: downstream tasks may be inaccessible, costly to run, or unavailable at compression time. These limitations highlight an urgent and underexplored problem: Compressed Feature Quality Assessment (CFQA) — how can we estimate the semantic quality of compressed features without relying on downstream inference?
Solving CFQA presents several challenges. First, there is no standard dataset that provides compressed features with corresponding task performance across multiple tasks and codecs. Second, existing similarity metrics lack validation for high-dimensional features and often fail to generalize across tasks. Third, there is no unified protocol for evaluating whether a given metric faithfully reflects semantic degradation.
To bridge this gap, we take the first step toward systematic study of CFQA. In contrast to prior works that rely on task supervision or task-specific codecs, we treat CFQA as a standalone problem and aim to benchmark its core components: datasets, metrics, and evaluation protocols.
We make the following contributions.
  • Multi-Task Benchmark Dataset: We construct the first CFQA dataset, consisting of 300 original features and 12,000 compressed features from three vision tasks (classification, segmentation, depth estimation) and four representative codecs (including handcrafted and learned models). The dataset enables quantitative analysis of semantic degradation across tasks, bitrates, and codecs.
  • Truth Semantic Distortion Derivation: For each compressed feature, we provide task-specific semantic distortion labels (rank change, mIoU drop, RMSE difference) computed by comparing task head outputs using original and compressed features. These labels serve as the ground truth for training and evaluating quality metrics.
  • Existing Metric Evaluation: We evaluate three representative signal-based metrics(MSE, cosine similarity, and Centered Kernel Alignment (CKA)) and analyze their sensitivity to compression and their correlation with task-specific degradation. Our analysis reveals their varying effectiveness, stability, and failure cases, providing critical insights into their applicability for semantic quality prediction.
We hope this work lays the foundation for future CFQA research and bridges the gap between low-level compression and high-level semantic utility. By providing a standardized benchmark and analysis pipeline, our work encourages the development of lightweight, generalizable, and task-agnostic quality metrics that support real-world deployment of compressed features.
2. Related Work
2.1. Feature Coding
Feature coding has received growing attention (Zhang et al., 2021; Kim et al., 2023; Misra et al., 2022; Alvar and Bajić, 2019, 2020; Guo et al., 2023; WG2, 2023; Yan et al., 2021; Choi and Bajić, 2018; Chen et al., 2020b, a; Gao et al., 2024a; Zhu et al., 2024; Gao et al., 2025; Ma et al., 2024) in the cloud-edge collaborative intelligence scenarios at early stage. Recently, large model feature coding has attracted increasing interest (Oquab et al., 2023; Guo et al., 2025). These work extend feature coding to the scenarios where features are transmitted, stored, and reused.
2.2. Semantic Distortion Measurement
Semantic distortion measurement methods are classified into three main categories: signal fidelity metrics, semantics fidelity metrics, and task-based metrics. Signal fidelity metrics (Choi and Bajić, 2018; Cai et al., 2022; Wang et al., 2022; Niu et al., 2021; Kim et al., 2023) focus on measuring the distortion between original and reconstructed features using traditional metrics like MSE. This approach is directly borrowed from image compression techniques, where the signal similarity is regarded as the compression distortion.
Semantics fidelity metrics (Alvar and Bajić, 2021, 2020; Choi and Bajić, 2021; Ikusan and Dai, 2021; Feng et al., 2022; Chen et al., 2019; Hu et al., 2020; Chen et al., 2020a; Duan et al., 2020) assess the preservation of semantic information, where distortion is measured based on the performance drop in specific machine vision tasks. These methods provide a more task-relevant measure of distortion, linking reconstruction quality directly to task performance.
Task-based metrics (Singh et al., 2020; Yang et al., 2024; Gao et al., 2024b; Zhang et al., 2021; Yan et al., 2021; Alvar and Bajić, 2019; Misra et al., 2022; Henzel et al., 2022; Wang et al., 2023; Li et al., 2023; Yuan et al., 2022) directly measure semantic distortion using the task performance. These methods are specialized for a particular task and lack generalizability.
Task Source Num. of Org. Feat. Propocessing Feature Codecs Num of Comp. Feat. Feature Shape GT Distortion
Cls ImageNet 100 Resize HM, VTM Multi-task Hyperprior Task-specific Hyperprior 4000 257$\times$1536 Rank
Seg VOC 2012 100 Flip and Crop 4000 2$\times$1370$\times$1536 mIoU difference
Dpt NYUv2 100 Flip 4000 2$\times$4$\times$161$\times$1536 RMSE difference
Table 1. Abstract information of the proposed dataset for CFQA. (Refer to Sec. 4 for more details.)
3. Problem Formulation and Applications
3.1. Problem Formulation
Given a pretrained model $\mathcal{M}$, we denote its extracted intermediate feature from an input $x\in\mathcal{X}$ as $\mathbf{f}=\mathcal{M}(x)\in\mathbb{R}^{d}$, where $d$ is the feature dimension. A feature codec $\mathcal{C}$ encodes $\mathbf{f}$ into a compact bitstream and decodes a compressed representation $\hat{\mathbf{f}}=\mathcal{C}^{-1}(\mathcal{C}(\mathbf{f}))$. Our objective is to evaluate how much semantic information is preserved in $\hat{\mathbf{f}}$ with respect to $\mathbf{f}$, without accessing downstream task ground-truth or executing inference.
We define this evaluation task as Feature Quality Assessment (FQA): Given a pair of original and compressed features $(\mathbf{f},\hat{\mathbf{f}})$, estimate the semantic quality $q\in\mathbb{R}$ of the compressed feature, such that $q$ strongly correlates with its performance on downstream tasks.
Assume a downstream task $\mathcal{T}$ with a head network $h_{\mathcal{T}}$, producing an output $y=h_{\mathcal{T}}(\cdot)$. Let the task-specific performance metric be denoted as $\mathcal{A}_{\mathcal{T}}(\cdot)$. When using compressed features, the task performance becomes: $s=\mathcal{A}_{\mathcal{T}}(h_{\mathcal{T}}(\hat{\mathbf{f}}))$. This value $s\in\mathbb{R}$ reflects the ground-truth semantic utility of the compressed feature. However, computing $s$ requires task execution and labels, which may be unavailable or expensive. The goal of FQA is to estimate a score $q=Q(\mathbf{f},\hat{\mathbf{f}})$, where $Q(\cdot)$ is a task-agnostic quality metric, such that $q\approx s$ in terms of correlation across samples.
3.2. Application Scenarios
We illustrate two typical application scenarios of CFQA in Fig. 1.
First, CFQA can be directly integrated into the codec training process as a supervisory signal to align compression objectives with downstream task performance. Recent studies (Singh et al., 2020; Gao et al., 2024a) show that task-aware feature coding benefits significantly from semantic-aware supervision. CFQA offers such guidance without requiring end-to-end task labels. As shown on the left of Fig. 1, both the original and reconstructed features are input to the CFQA module, which estimates the semantic distortion. The distortion score is then used to generate gradients for optimizing the encoder and decoder. With a well-designed CFQA metric, the resulting codec learns to preserve semantic content even under low-bitrate constraints.
Second, CFQA is essential in edge-cloud collaborative systems, where features are extracted at the edge and transmitted to the cloud for inference. Since the downstream model resides on the cloud side, semantic distortion cannot be directly measured at the edge. Here, CFQA is used as a proxy to estimate task-relevant degradation. As illustrated on the right of Fig. 1, the compressed feature is first evaluated by the CFQA module. If it is judged to be of high quality, the bitstream is transmitted to the cloud. Otherwise, the edge device re-encodes the feature at a higher bitrate to better preserve semantic information. This strategy ensures the reliability of transmitted features while reducing unnecessary bandwidth consumption.
Although our work focuses on semantic degradation due to compression, we emphasize that the value of CFQA extends beyond codec benchmarking: it serves as a crucial component in a variety of systems where features are extracted, compressed, transmitted, cached, or reused.
4. Dataset Construction
4.1. Overview
Our proposed dataset is designed to support the analysis of semantic distortion introduced by lossy compression. It includes three tasks, 300 original features, and 12000 compressed features from 4 kinds of feature codecs. The feature cover diverse image processing methods and multiple feature extraction strategies. The abstract information of the dataset is presented in Table 1.
4.2. Model and Task Selection
Since feature coding still focus on visual signals, we start CFQA by visual features. We choose DINOv2 (Oquab et al., 2023) as the backbone model for feature extraction due to its strong generalization ability and wide adoption in general-purpose vision tasks. To cover a broad range of semantic complexity, we select three widely studied vision tasks: image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt). These tasks span coarse-to-fine semantic understanding: Cls focuses on image-level category prediction, Seg introduces spatial semantics with class-wise alignment at the pixel level, and Dpt requires detailed geometry prediction.
By including these three tasks, we aim to ensure that semantic distortion is evaluated across multiple levels of abstraction, which is essential for assessing the general applicability and sensitivity of CFQA metrics.
4.3. Source Data Collection
To ensure semantic diversity, we select 100 representative samples for each task. For Cls, we sample 100 ImageNet (Deng et al., 2009) images from 100 distinct categories, each correctly predicted by the DINOv2 classifier. For Seg, we sample 100 images from the Pascal VOC 2012 (Everingham et al., 2010) validation set, covering all 20 semantic categories. For Dpt, we sample 100 images from the NYUv2 (Nathan Silberman and Fergus, 2012) dataset, spanning all 16 scenes.
This sampling ensures feature diversity while maintaining dataset size suitable for controlled evaluation.
4.4. Original Feature Collection
[…]
The various image preprocessing and split points simulate real-world input variation and improve generalizability.
4.5. Compressed Feature Collection
To simulate different types and strengths of semantic degradation, we compress original features through four codecs. All original features are flattened into 2D arrays before encoding, and compressed features are reshaped to original formats after decoding.
Handcrafted Codecs
We uniformly quantize original features to [0, 1023], followed by two handcrafted codecs: HM Intra coding (configured with encoder_intra_main_rext.cfg) and VTM Intra coding: (configured with encoder_intra_vtm.cfg). For both codec, we use the YUV-400 format for feature coding and the quantization parameters set to $\in{2,4,6,\ldots,20}$ to simulate different bitrate levels and distortion strengths.
Learning-Based Codecs
We adopt the Hyperprior model (Ballé et al., 2018), which serves as the milestone for learned feature coding. Hyperprior models offer a good balance between modeling capacity and computational efficiency. Most existing compression methods follow the Hyperprior architecture. Therefore, it can simulate the semantic distortion introduced by learning-based methods.
To analyze optimization effects, we train two variants: Task-Specific: The codec is trained independently on features extracted from a specific task. Multi-Task: The codec is trained on features extracted from the three tasks. All codecs are trained using the training strategies proposed in (Gao et al., 2024c), and the lambda information can be found in our released models.
4.6. True Semantic Distortion Collection
To evaluate the effectiveness of CFQA metrics, we need to compare them with true semantic distortion. We define the true semantic distortion distortion as the task performance difference between the compressed featrure $\hat{\mathbf{f}}$ and the original feature $\mathbf{f}$.
For Cls, we compute the rank of the ground-truth label in the prediction logits derived from $\hat{\mathbf{f}}$. Higher rank indicates greater semantic degradation. The rank value of the original feature is 1. For Seg, we compute the mIoU difference between the segmentation masks predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$. For Dpt, we compute the RMSE difference between the depth maps predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$.
These task-specific scores serve as ground-truth semantic distortion for evaluating the effectiveness of feature quality metrics.
HM VTM
Cls Seg Dpt Cls Seg Dpt
BPFP Acc. BPFP mIoU BPFP RMSE BPFP Acc. BPFP mIoU BPFP RMSE
32 100 32 83.39 32 0.37 32 100 32 83.39 32 0.37
0.006 0.00 0.002 4.59 0.19 1.73 0.006 1.00 0.004 22.56 0.23 1.59
0.008 0.00 0.004 11.39 0.37 1.40 0.01 1.00 0.01 40.18 0.36 1.36
0.012 2.00 0.009 32.22 0.64 1.27 0.02 10.00 0.02 47.59 0.52 1.25
0.02 6.00 0.02 47.79 1.01 1.00 0.03 13.00 0.04 55.05 0.75 1.23
0.04 18.00 0.05 59.63 1.41 0.81 0.06 26.00 0.09 65.24 1.02 1.06
0.08 25.00 0.11 67.87 1.82 0.63 0.11 44.00 0.17 72.34 1.26 0.92
0.15 55.00 0.21 74.54 2.16 0.55 0.23 81.00 0.31 76.81 1.52 0.83
0.33 84.00 0.41 78.40 2.55 0.47 0.49 92.00 0.56 79.79 1.85 0.73
0.67 94.00 0.73 80.49 2.91 0.43 0.86 97.00 0.91 81.33 2.23 0.59
1.12 97.00 1.13 81.86 3.28 0.40 1.26 98.00 1.28 82.04 2.67 0.52
Table 2. Rate-accuracy performance of handcrafted codecs, with the first row showing original features’ performance.
Multi-Task-Trained Hyperprior Task-Specific-Trained Hyperprior
Cls Seg Dpt Cls Seg Dpt
BPFP Acc. BPFP mIoU BPFP RMSE BPFP Acc. BPFP mIoU BPFP RMSE
32 100 32 83.39 32 0.37 32 100 32 83.39 32 0.37
0.01 0.00 0.0001 2.65 0.0003 1.68 0.34 16.00 0.05 50.67 0.12 2.09
0.12 18.00 0.10 58.74 0.10 1.79 0.43 35.00 0.08 60.69 0.18 1.66
0.58 44.00 0.46 76.70 0.52 1.55 0.56 55.00 0.14 65.45 0.30 1.22
0.71 52.00 0.57 77.80 0.64 1.24 0.65 61.00 0.21 71.89 0.42 0.88
1.01 59.00 0.85 79.30 0.91 0.99 1.15 70.00 0.29 74.76 0.57 0.82
1.19 67.00 1.01 80.06 1.07 1.45 1.39 72.00 0.42 78.10 0.70 0.73
1.34 77.00 1.14 80.50 1.20 0.77 1.81 78.00 0.65 79.09 1.06 0.60
1.45 81.00 1.23 80.66 1.29 0.92 1.96 81.00 0.94 79.59 1.30 0.49
1.75 94.00 1.45 81.22 1.52 0.76 2.15 84.00 1.41 81.02 1.39 0.44
2.18 96.00 1.77 81.73 1.84 0.42 2.36 88.00 1.65 82.02 1.48 0.43
Table 3. Rate-accuracy performance of learning-based codecs, with the first row showing original features’ performance.
MSE Cosine Similarity CKA
Codec Task PLCC SROCC PLCC SROCC PLCC SROCC
HM Cls 0.6641 0.9113 -0.7368 -0.9116 -0.5788 -0.9089
Seg -0.6210 -0.6729 0.6866 0.6729 0.6133 0.6726
Dpt 0.8601 0.9036 -0.8799 -0.9036 -0.8829 -0.9027
VTM Cls 0.6220 0.8939 -0.7089 -0.8939 -0.5499 -0.8932
Seg -0.4784 -0.5213 0.5131 0.5213 0.4885 0.5213
Dpt 0.8344 0.8750 -0.8572 -0.8747 -0.8478 -0.8754
Hyperprior (Multi-Task) Cls -0.0281 0.4707 -0.8907 -0.7165 -0.3277 -0.6593
Seg 0.0595 -0.0251 0.3496 0.0604 -0.0059 0.0847
Dpt 0.6466 0.6675 -0.4752 -0.6528 -0.5876 -0.6004
Hyperprior (Task-Specific) Cls -0.0220 0.1303 -0.5486 -0.6084 -0.2622 -0.3365
Seg 0.3259 0.3235 -0.0552 -0.1669 -0.3138 -0.2187
Dpt -0.0787 0.4915 -0.8427 -0.8258 -0.5483 -0.5638
Table 4. Average CFQA performance of the three baseline metrics on Cls, Seg, and Dpt tasks.
5. Experiments and Analysis
5.1. Baseline CFQA Metrics
To cover diverse types of distortions and similarity relationships between features, we select three representative signal-based metrics: MSE, Cosine Similarity, and Centered Kernel Alignment (CKA)  (Kornblith et al., 2019).
  • MSE: MSE measures feature distortion at the element level. It is widely used in signal processing fields.
  • Cosine Similarity: Cosine similarity measures the angular difference between feature vectors, normalized by their magnitudes. This metric captures directional alignment in the feature space, which is often more robust to scale distortions and is considered to better reflect semantic similarity in high-dimensional embeddings (Radford et al., 2021).
  • CKA: CKA measures similarity between two features using normalized HSIC (Hilbert-Schmidt Independence Criterion). Unlike MSE or cosine similarity, CKA captures structural relationships and global correlation patterns between two features, making it suitable for comparing multi-token or multi-patch feature distributions as commonly found in vision transformers.
These metrics form diverse baselines from element-level to structure-level distortion metrics.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2. PLCC distribution visualization of MSE, cosine similarity, and CKA. The first, second, and third rows correspond to the Cls, Seg, and Dpt tasks, respectively. The first, second, third, and fourth column correspond to the HM, VTM, multi-task-trained Hyperprior, and task-specific-trained Hyperprior codecs.
5.2. Evaluation Protocol
We define a evaluation protocol to assess how well a CFQA metric predicts the semantic distortion of compressed features. PLCC and SROCC are used in our experiments.
  • PLCC (Linearity): PLCC quantifies how well the predicted quality scores approximate the actual semantic distortion in a linear sense.
  • SROCC (Monotonicity): SROCC assesses the consistency in the ranking between predicted and true distortions.
Specifically, for each original feature $\mathbf{f}_{i}$, we compute these two metrics on its 10 predicted quality scores and 10 true semantic distortion. Each metric is evaluated on all three tasks and across all codecs.
5.3. Rate-Accuracy Performance Analysis
Table 2 presents the rate-accuracy performance of the handcrafted codecs. For all three tasks, both codecs exhibit a broad range of bitrates and corresponding accuracy levels, spanning from nearly lossless reconstruction to significant semantic degradation. This wide span underscores their utility in effectively simulating traditional coding distortions. However, the Dpt task shows a narrower variation in both bitrate and performance. This can be attributed to the use of multi-scale features, which include lower-level layers with higher redundancy and reduced semantic abstraction, resulting in less fluctuation in both bitrate and distortion.
Table 3 shows the rate-accuracy performance of two learning-based codecs. In contrast to handcrafted codecs, the learning-based codecs exhibit more varied performance behaviors across tasks. Specifically, the multi-task codec outperforms the task-specific codec in Cls, whereas the task-specific codec demonstrates superior performance in Dpt. This suggests that incorporating fine-grained task features enhances the compression training of coarse-grained features, while fine-grained features are more effectively trained using a single distribution tailored to the task.
These four codecs comprehensively simulate a wide range of compression distortions, making them not only highly suitable but also essential for supporting research in compressed feature quality assessment.
5.4. CFQA Performance Analysis
5.4.1. Average Performance Analysis
Table 4 presents the average CFQA performance of the three baseline metrics. Overall, the three baseline metrics show higher PLCC and SROCC values for the handcrafted codecs compared to the learning-based codecs. This indicates that the semantic distortions introduced by handcrafted codecs are more stable. The larger fluctuations in semantic distortion observed for learning-based codecs are likely due to the inherent variability in the training process, as these models are data-driven and task-specific. Additionally, since handcrafted codecs are block-based, they tend to exhibit more consistent and predictable distortion patterns.
Among the three baseline metrics, cosine similarity consistently demonstrates a higher degree of linearity and monotonicity in relation to true semantic distortion. This is due to the metric’s ability to capture angular relationships between feature vectors, making it particularly effective for high-dimensional feature sets such as those produced by ViT.
For handcrafted codecs, the three baseline metrics show better performance in capturing HM-generated distortions compared to VTM. This is likely because VTM employs more complex coding tools, resulting in more intricate and potentially less predictable distortion patterns. In contrast, for learning-based codecs, significant performance variation is observed across the three baseline metrics. This variability is expected, as learning-based encoders produce a wider range of distortion types, making it more challenging for simple metrics to achieve linear fitting.
Among the three tasks, Seg presents the most complex and difficult-to-fit distortions. For most codecs, the baseline metrics show poorer fitting for Seg compared to Cls and Dpt. This reflects the higher semantic complexity in Seg, which is more challenging for these metrics to capture accurately.
5.4.2. Distribution Analysis
Figure 2 visualizes the distribution of PLCC. To provide a clearer view, all PLCC values are rounded to the nearest tenth before computing the frequency histograms.
Overall, the PLCC distributions for handcrafted codecs are more concentrated compared to those of learning-based codecs. This aligns with the fact that handcrafted codecs introduce more consistent and predictable semantic distortion patterns.
Among the three baseline metrics, cosine similarity exhibits the most concentrated PLCC distributions. This observation aligns with its superior average PLCC values, as reported in Table 2 and Table 3, indicating that cosine similarity provides a more stable quality estimation across varying conditions.
Across all three metrics, Seg shows more dispersed PLCC distributions compared to Cls and Dpt. In some cases, such as with CKA, both positive and negative correlations are observed within the same metric. This highlights the high complexity of distortion patterns in segmentation features. The broader distribution further confirms that existing baseline metrics struggle to model such complex semantic degradation accurately.
5.5. Limitations
While the evaluated metrics demonstrate moderate performance in certain tasks and codecs, they also exhibit significant limitations. First, signal-based metrics like MSE lack task sensitivity and often fail to reflect meaningful semantic degradation, especially in segmentation and generation tasks. Second, cosine and CKA metrics, although capturing structural alignment, are sometimes unstable under nonlinear distortions introduced by learned codecs. Finally, none of the metrics generalize consistently across all tasks and compression strategies, suggesting that handcrafted similarity alone may not suffice for universal FQA. These limitations highlight the need for more adaptive, task-aware, or learned semantic quality estimators, which we consider for future work.
6. Conclusion
This paper introduces the concept of Compressed Feature Quality Assessment (CFQA), a crucial area for evaluating the semantic degradation of compressed features in systems where features are transmitted, stored, and reused. We present the first benchmark dataset for CFQA, which lays the groundwork for further research in this field. We assess the widely used metric in CFQA and provide insights for interested researchers. Future research will focus on developing adaptive CFQA metrics that generalize across tasks and feature coding strategies.

### v2

@@
 In this context, feature coding, which compresses features into compact bitstreams, becomes a critical component for scenarios involving feature transmission, storage, and reuse.
-However, this compression process introduces inherent semantic degradation that is notoriously difficult to quantify with traditional metrics.
-To address this, this paper introduces the research problem of Compressed Feature Quality Assessment (CFQA), which seeks to evaluate the semantic fidelity of compressed features.
+However, this compression process inevitably introduces semantic degradation that is difficult to quantify with traditional metrics.
+To address this, we formalize the research problem of Compressed Feature Quality Assessment (CFQA), aiming to evaluate the semantic fidelity of compressed features.
 To advance CFQA research, we propose the first benchmark dataset, comprising 300 original features and 12000 compressed features derived from three vision tasks and four feature codecs.
-Task-specific performance drops are provided as true semantic distortion for the evaluation of CFQA metrics.
-We assess the performance of three widely used metrics (MSE, cosine similarity, and Centered Kernel Alignment) in capturing semantic degradation.
-The results underscore the representativeness of the dataset and highlight the need for more refined metrics capable of addressing the nuances of semantic distortion in compressed features.
-To facilitate the ongoing development of CFQA research, we release the dataset and all accompanying source code at https://github.com/chansongoal/Compressed-Feature-Quality-Assessment.
-This contribution aims to advance the field and provide a foundational resource for the community to explore CFQA.
-Compressed Feature Quality Assessment, Coding for Machines, Feature Coding
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Multimedia databasesccs: Computing methodologies Image compressionccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
+Task-specific performance degradation is provided as true semantic distortion for evaluating CFQA metrics.
+We systematically assess three widely used metrics – MSE, cosine similarity, and Centered Kernel Alignment (CKA) – in terms of their ability to capture semantic degradation.
+Our findings demonstrate the representativeness of the proposed dataset while underscoring the need for more sophisticated metrics capable of measuring semantic distortion in compressed features.
+This work advances the field by establishing a foundational benchmark and providing a critical resource for the community to explore CFQA.
+To foster further research, we release the dataset and all associated source code at https://github.com/chansongoal/Compressed-Feature-Quality-Assessment.
+Compressed Feature Quality Assessment (CFQA), Coding for Machines, Feature Coding
+journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758309isbn: 979-8-4007-2035-2/2025/10ccs: Information systems Multimedia databasesccs: Computing methodologies Image compressionccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
 Figure 1. Exemplar application scenarios of compressed feature quality assessment.
 1. Introduction
-The rapid deployment of large foundation models (e.g., DINOv2, LLaMA3) across distributed and resource-constrained environments has created a growing need to transmit intermediate features rather than raw signals.
+Rapid deployment of large foundation models (e.g., DINOv2 (Oquab et al., 2023), LLaMA3 (Dubey et al., 2024)) in distributed and resource-constrained environments has created a growing need to transmit intermediate features rather than raw signals (Gao et al., 2024c).
 Feature coding, which compresses these intermediate representations, plays a vital role in enabling scalable, privacy-preserving, and efficient systems.
-Unlike traditional image or video coding that prioritizes perceptual quality, feature coding targets the preservation of task-relevant semantics embedded in deep representations.
-However, feature coding inevitably introduces semantic degradation: a loss of discriminative information that can compromise downstream performance.
-This degradation is fundamentally different from pixel-level distortions and often cannot be captured by conventional distortion metrics such as MSE or PSNR.
-While task accuracy is a more reliable indicator of semantic quality, it is impractical for deployment: downstream tasks may be inaccessible, costly to run, or unavailable at compression time.
-These limitations highlight an urgent and underexplored problem: Compressed Feature Quality Assessment (CFQA) — how can we estimate the semantic quality of compressed features without relying on downstream inference?
+Unlike traditional image or video coding that prioritizes perceptual quality, feature coding targets the preservation of task-relevant semantics embedded in feature representations.
+However, feature coding inevitably introduces semantic degradation: a loss of semantic information that may compromise the performance of large models.
+This degradation is fundamentally different from pixel-level distortions and often cannot be captured by conventional distortion metrics such as MSE or SSIM.
+Although task accuracy is a more reliable measurement of semantic distortion, it is impractical for practical deployments: downstream tasks may be inaccessible, costly to run, or unavailable in the feature coding process.
+These limitations highlight an urgent and underexplored problem: Compressed Feature Quality Assessment (CFQA) – How can we estimate the semantic distortion of compressed features without relying on downstream inference?
 Solving CFQA presents several challenges.
-First, there is no standard dataset that provides compressed features with corresponding task performance across multiple tasks and codecs.
-Second, existing similarity metrics lack validation for high-dimensional features and often fail to generalize across tasks.
-Third, there is no unified protocol for evaluating whether a given metric faithfully reflects semantic degradation.
-To bridge this gap, we take the first step toward systematic study of CFQA.
-In contrast to prior works that rely on task supervision or task-specific codecs, we treat CFQA as a standalone problem and aim to benchmark its core components: datasets, metrics, and evaluation protocols.
-We make the following contributions.
-  • Multi-Task Benchmark Dataset: We construct the first CFQA dataset, consisting of 300 original features and 12,000 compressed features from three vision tasks (classification, segmentation, depth estimation) and four representative codecs (including handcrafted and learned models).
-The dataset enables quantitative analysis of semantic degradation across tasks, bitrates, and codecs.
-  • Truth Semantic Distortion Derivation: For each compressed feature, we provide task-specific semantic distortion labels (rank change, mIoU drop, RMSE difference) computed by comparing task head outputs using original and compressed features.
-These labels serve as the ground truth for training and evaluating quality metrics.
-  • Existing Metric Evaluation: We evaluate three representative signal-based metrics(MSE, cosine similarity, and Centered Kernel Alignment (CKA)) and analyze their sensitivity to compression and their correlation with task-specific degradation.
-Our analysis reveals their varying effectiveness, stability, and failure cases, providing critical insights into their applicability for semantic quality prediction.
-We hope this work lays the foundation for future CFQA research and bridges the gap between low-level compression and high-level semantic utility.
-By providing a standardized benchmark and analysis pipeline, our work encourages the development of lightweight, generalizable, and task-agnostic quality metrics that support real-world deployment of compressed features.
+First, there is no public dataset that provides compressed features with corresponding task performance across multiple tasks and codecs.
+Second, existing similarity metrics lack validation for high-dimensional features and often fail to generalize across various tasks.
+Third, the field lacks a unified evaluation protocol to assess whether a given metric reliably measures semantic distortion.
+To bridge this gap, we take the first step towards a systematic study of CFQA.
+In contrast to previous works that rely on task supervision or task-specific codecs, we treat CFQA as a standalone problem and aim to benchmark its core components: dataset, metrics, and evaluation protocols.
+In summary, we make the following contributions:
+  • Benchmark Dataset: We construct the first comprehensive CFQA dataset, consisting of 300 original features and 12000 compressed features from three vision tasks (image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt)) and four representative codecs (handcrafted and learning-based).
+The dataset enables quantitative analysis of semantic distortion across tasks, bitrates, and coding methods.
+  • Ground-truth Semantic Distortion: For each compressed feature, we provide task-specific semantic distortion labels computed by comparing task head outputs using original and compressed features.
+These labels serve as the ground truth for training and evaluating semantic distortion metrics.
+  • Baseline Metric Evaluation: We evaluate three representative metrics (MSE, cosine similarity, and Centered Kernel Alignment (CKA)) and analyze their sensitivity to feature coding and their ability to measure task-specific semantic distortion.
+Our analysis reveals their varying strengths and limitations, providing insight into their applicability to semantic distortion measurement.
+By establishing a standardized benchmark and evaluation framework, this work bridges the gap between low-level compression and high-level semantic utility.
+We hope to catalyze the development of lightweight, generalizable, and task-agnostic CFQA metrics that advance the real-world adoption of feature coding.
+Our dataset and code are publicly available to foster reproducibility and future research.
 2. Related Work
 2.1. Feature Coding
-Feature coding has received growing attention (Zhang et al., 2021; Kim et al., 2023; Misra et al., 2022; Alvar and Bajić, 2019, 2020; Guo et al., 2023; WG2, 2023; Yan et al., 2021; Choi and Bajić, 2018; Chen et al., 2020b, a; Gao et al., 2024a; Zhu et al., 2024; Gao et al., 2025; Ma et al., 2024) in the cloud-edge collaborative intelligence scenarios at early stage.
-Recently, large model feature coding has attracted increasing interest (Oquab et al., 2023; Guo et al., 2025).
-These work extend feature coding to the scenarios where features are transmitted, stored, and reused.
+Feature coding has received growing attention (Zhang et al., 2021; Kim et al., 2023; Misra et al., 2022; Alvar and Bajić, 2019, 2020; Guo et al., 2023; WG2, 2023; Yan et al., 2021; Choi and Bajić, 2018; Chen et al., 2020b, a; Gao et al., 2024a; Zhu et al., 2024; Gao et al., 2025a; Ma et al., 2024; Chen et al., 2024) in edge-cloud collaborative intelligence scenarios.
+Recently, large model feature coding has attracted increasing interest (Gao et al., 2024c, 2025b, 2025c).
+These works extend feature coding to scenarios where features are transmitted, stored, and reused.
 2.2. Semantic Distortion Measurement
@@
 These methods are specialized for a particular task and lack generalizability.
-Task Source Num. of Org. Feat. Propocessing Feature Codecs Num of Comp. Feat. Feature Shape GT Distortion
-Cls ImageNet 100 Resize HM, VTM Multi-task Hyperprior Task-specific Hyperprior 4000 257$\times$1536 Rank
-Seg VOC 2012 100 Flip and Crop 4000 2$\times$1370$\times$1536 mIoU difference
-Dpt NYUv2 100 Flip 4000 2$\times$4$\times$161$\times$1536 RMSE difference
+Task Source Num. of Org. Feat. Pre-processing Feature Codecs Num. of Comp. Feat. Feature Shape GT Distortion
+Cls ImageNet 100 Resize
+HM, VTM
+Multi-task Hyperprior
+Task-specific Hyperprior
+[…]
+[…]
+[…]
 Table 1. Abstract information of the proposed dataset for CFQA. (Refer to Sec. 4 for more details.)
@@
 Given a pretrained model $\mathcal{M}$, we denote its extracted intermediate feature from an input $x\in\mathcal{X}$ as $\mathbf{f}=\mathcal{M}(x)\in\mathbb{R}^{d}$, where $d$ is the feature dimension.
-A feature codec $\mathcal{C}$ encodes $\mathbf{f}$ into a compact bitstream and decodes a compressed representation $\hat{\mathbf{f}}=\mathcal{C}^{-1}(\mathcal{C}(\mathbf{f}))$.
+A feature codec $\mathcal{C}$ encodes $\mathbf{f}$ into a compact bitstream and decodes it back to $\hat{\mathbf{f}}=\mathcal{C}^{-1}(\mathcal{C}(\mathbf{f}))$.
 Our objective is to evaluate how much semantic information is preserved in $\hat{\mathbf{f}}$ with respect to $\mathbf{f}$, without accessing downstream task ground-truth or executing inference.
-We define this evaluation task as Feature Quality Assessment (FQA): Given a pair of original and compressed features $(\mathbf{f},\hat{\mathbf{f}})$, estimate the semantic quality $q\in\mathbb{R}$ of the compressed feature, such that $q$ strongly correlates with its performance on downstream tasks.
-Assume a downstream task $\mathcal{T}$ with a head network $h_{\mathcal{T}}$, producing an output $y=h_{\mathcal{T}}(\cdot)$.
-Let the task-specific performance metric be denoted as $\mathcal{A}_{\mathcal{T}}(\cdot)$.
-When using compressed features, the task performance becomes: $s=\mathcal{A}_{\mathcal{T}}(h_{\mathcal{T}}(\hat{\mathbf{f}}))$.
-This value $s\in\mathbb{R}$ reflects the ground-truth semantic utility of the compressed feature.
-However, computing $s$ requires task execution and labels, which may be unavailable or expensive.
-The goal of FQA is to estimate a score $q=Q(\mathbf{f},\hat{\mathbf{f}})$, where $Q(\cdot)$ is a task-agnostic quality metric, such that $q\approx s$ in terms of correlation across samples.
+[…]
+[…]
 3.2. Application Scenarios
-We illustrate two typical application scenarios of CFQA in Fig. 1.
-First, CFQA can be directly integrated into the codec training process as a supervisory signal to align compression objectives with downstream task performance.
+We illustrate two typical application scenarios of CFQA in Figure 1.
+In the first scenario (left of Figure 1), CFQA can be directly integrated into the codec training process as a supervisory signal to align compression objectives with downstream task performance.
 Recent studies (Singh et al., 2020; Gao et al., 2024a) show that task-aware feature coding benefits significantly from semantic-aware supervision.
 CFQA offers such guidance without requiring end-to-end task labels.
-As shown on the left of Fig.
-1, both the original and reconstructed features are input to the CFQA module, which estimates the semantic distortion.
+As shown on the left of Figure 1, both the original and reconstructed features are input to the CFQA module, which estimates the semantic distortion.
 The distortion score is then used to generate gradients for optimizing the encoder and decoder.
-With a well-designed CFQA metric, the resulting codec learns to preserve semantic content even under low-bitrate constraints.
-Second, CFQA is essential in edge-cloud collaborative systems, where features are extracted at the edge and transmitted to the cloud for inference.
+With a well-designed CFQA metric, the resulting codec learns to preserve semantic information even under low-bitrate constraints.
+In the second scenario (right of Figure 1), CFQA is essential in edge-cloud collaborative systems, where features are extracted at the edge and transmitted to the cloud for inference.
 Since the downstream model resides on the cloud side, semantic distortion cannot be directly measured at the edge.
-Here, CFQA is used as a proxy to estimate task-relevant degradation.
-As illustrated on the right of Fig.
-1, the compressed feature is first evaluated by the CFQA module.
+Here, CFQA is used as a proxy to estimate task-relevant semantic distortion.
+As illustrated on the right of Figure 1, the compressed feature is first evaluated by the CFQA module.
 If it is judged to be of high quality, the bitstream is transmitted to the cloud.
@@
 This strategy ensures the reliability of transmitted features while reducing unnecessary bandwidth consumption.
-Although our work focuses on semantic degradation due to compression, we emphasize that the value of CFQA extends beyond codec benchmarking: it serves as a crucial component in a variety of systems where features are extracted, compressed, transmitted, cached, or reused.
+Although our work focuses on semantic distortion due to compression, we emphasize that the value of CFQA extends beyond codec benchmarking: it serves as a crucial component in a variety of systems where features are extracted, compressed, transmitted, cached, or reused.
 4. Dataset Construction
@@
 Our proposed dataset is designed to support the analysis of semantic distortion introduced by lossy compression.
-It includes three tasks, 300 original features, and 12000 compressed features from 4 kinds of feature codecs.
-The feature cover diverse image processing methods and multiple feature extraction strategies.
-The abstract information of the dataset is presented in Table 1.
+It includes three tasks, 300 original features, and 12000 compressed features from 4 feature codecs.
+The features cover diverse image processing methods and multiple feature extraction strategies.
+The overall information of the dataset is presented in Table 1.
 4.2. Model and Task Selection
-Since feature coding still focus on visual signals, we start CFQA by visual features.
-We choose DINOv2 (Oquab et al., 2023) as the backbone model for feature extraction due to its strong generalization ability and wide adoption in general-purpose vision tasks.
-To cover a broad range of semantic complexity, we select three widely studied vision tasks: image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt).
+Since current feature coding research primarily focuses on visual signals, we initiate the study of CFQA with visual features.
+We adopt DINOv2 (Oquab et al., 2023) as our backbone feature extractor due to its strong generalization capability and widespread adoption in general-purpose vision tasks.
+To cover a broad range of semantic distortion, we select three widely studied vision tasks: image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt).
 These tasks span coarse-to-fine semantic understanding: Cls focuses on image-level category prediction, Seg introduces spatial semantics with class-wise alignment at the pixel level, and Dpt requires detailed geometry prediction.
-By including these three tasks, we aim to ensure that semantic distortion is evaluated across multiple levels of abstraction, which is essential for assessing the general applicability and sensitivity of CFQA metrics.
+This multi-task approach is critical for evaluating the generalizability and sensitivity of CFQA metrics, ensuring they remain robust across different semantic requirements.
 4.3. Source Data Collection
@@
 For Dpt, we sample 100 images from the NYUv2 (Nathan Silberman and Fergus, 2012) dataset, spanning all 16 scenes.
-This sampling ensures feature diversity while maintaining dataset size suitable for controlled evaluation.
+This sampling strategy balances feature diversity with manageable dataset scale, enabling rigorous yet tractable evaluation.
 4.4. Original Feature Collection
 […]
-The various image preprocessing and split points simulate real-world input variation and improve generalizability.
+The inclusion of diverse image pre-processing techniques and varying split points emulates real-world input variability, enabling rigorous evaluation of CFQA metric generalizability.
 4.5. Compressed Feature Collection
-To simulate different types and strengths of semantic degradation, we compress original features through four codecs.
-All original features are flattened into 2D arrays before encoding, and compressed features are reshaped to original formats after decoding.
+To simulate different types and strengths of semantic distortion, we compress original features through four codecs. All original features are flattened into 2D arrays before encoding.
 Handcrafted Codecs
-We uniformly quantize original features to [0, 1023], followed by two handcrafted codecs: HM Intra coding (configured with encoder_intra_main_rext.cfg) and VTM Intra coding: (configured with encoder_intra_vtm.cfg).
-For both codec, we use the YUV-400 format for feature coding and the quantization parameters set to $\in{2,4,6,\ldots,20}$ to simulate different bitrate levels and distortion strengths.
+We select two handcrafted codecs: HM Intra coding (configured with encoder_intra_main_rext.cfg) and VTM Intra coding (configured with encoder_intra_vtm.cfg).
+Before encoding, we first uniformly quantize the original feature values to [0, 1023].
+For both codecs, we use the YUV 4:0:0 (monochrome) format for feature coding and set the quantization parameters to $\in\{2,4,6,8,\ldots,16,18,20\}$ to simulate various bitrates (measured by Bits Per Feature Point, BPFP) and distortion levels.
 Learning-Based Codecs
-We adopt the Hyperprior model (Ballé et al., 2018), which serves as the milestone for learned feature coding.
-Hyperprior models offer a good balance between modeling capacity and computational efficiency.
-Most existing compression methods follow the Hyperprior architecture.
-Therefore, it can simulate the semantic distortion introduced by learning-based methods.
-To analyze optimization effects, we train two variants: Task-Specific: The codec is trained independently on features extracted from a specific task.
-Multi-Task: The codec is trained on features extracted from the three tasks.
-All codecs are trained using the training strategies proposed in (Gao et al., 2024c), and the lambda information can be found in our released models.
-4.6. True Semantic Distortion Collection
-To evaluate the effectiveness of CFQA metrics, we need to compare them with true semantic distortion.
-We define the true semantic distortion distortion as the task performance difference between the compressed featrure $\hat{\mathbf{f}}$ and the original feature $\mathbf{f}$.
-For Cls, we compute the rank of the ground-truth label in the prediction logits derived from $\hat{\mathbf{f}}$.
-Higher rank indicates greater semantic degradation.
-The rank value of the original feature is 1.
+We adopt the Hyperprior model (Ballé et al., 2018) as it represents a milestone in learning-based feature coding.
+Since most learning-based feature coding methods build upon this architecture, it serves as an ideal testbed for studying the semantic distortion characteristics of learning-based methods.
+To investigate how optimization strategies affect compressed feature quality, we implement two distinct variants.
+Multi-Task Hyperprior Codec: The codec is trained on features extracted from the three tasks.
+Task-Specific Hyperprior Codec: The codec is trained exclusively on features extracted from a single task.
+All codecs are optimized following the training protocols established in (Gao et al., 2024c).
+The rate-distortion trade-off parameters ($\lambda$) can be found in our released models.
+4.6. Semantic Distortion Collection
+To establish a rigorous benchmark for evaluating CFQA metrics, we require ground-truth measurements of semantic distortion.
+We define the true semantic distortion as the performance degradation in downstream tasks when using compressed features $\hat{\mathbf{f}}$ versus original features $\mathbf{f}$.
+For Cls, we measure the deviation in prediction confidence by computing the rank in the softmax function generated from $\hat{\mathbf{f}}$.
+The original feature $\mathbf{f}$ achieves perfect ranking (rank=1), with higher ranks indicating more severe semantic degradation.
 For Seg, we compute the mIoU difference between the segmentation masks predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$.
 For Dpt, we compute the RMSE difference between the depth maps predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$.
-These task-specific scores serve as ground-truth semantic distortion for evaluating the effectiveness of feature quality metrics.
+These task-specific scores provide quantitative measures of semantic distortion, serving as the foundation for assessing CFQA metric performance.
 HM VTM
@@
 Table 2. Rate-accuracy performance of handcrafted codecs, with the first row showing original features’ performance.
-Multi-Task-Trained Hyperprior Task-Specific-Trained Hyperprior
+Multi-Task Hyperprior Task-Specific Hyperprior
 Cls Seg Dpt Cls Seg Dpt
@@
 Dpt 0.8344 0.8750 -0.8572 -0.8747 -0.8478 -0.8754
-Hyperprior (Multi-Task) Cls -0.0281 0.4707 -0.8907 -0.7165 -0.3277 -0.6593
+Hyperprior
+(Multi-Task)
+Cls -0.0281 0.4707 -0.8907 -0.7165 -0.3277 -0.6593
 Seg 0.0595 -0.0251 0.3496 0.0604 -0.0059 0.0847
 Dpt 0.6466 0.6675 -0.4752 -0.6528 -0.5876 -0.6004
-Hyperprior (Task-Specific) Cls -0.0220 0.1303 -0.5486 -0.6084 -0.2622 -0.3365
+Hyperprior
+(Task-Specific)
+Cls -0.0220 0.1303 -0.5486 -0.6084 -0.2622 -0.3365
 Seg 0.3259 0.3235 -0.0552 -0.1669 -0.3138 -0.2187
@@
 5.1. Baseline CFQA Metrics
-To cover diverse types of distortions and similarity relationships between features, we select three representative signal-based metrics: MSE, Cosine Similarity, and Centered Kernel Alignment (CKA)  (Kornblith et al., 2019).
+To cover diverse types of distortions and similarity relationships between features, we select three complementary metrics: MSE, cosine similarity, and CKA  (Kornblith et al., 2019).
+These metrics collectively span from local element-wise to global structural comparisons.
   • MSE: MSE measures feature distortion at the element level. It is widely used in signal processing fields.
   • Cosine similarity: Cosine similarity measures the angular difference between feature vectors, normalized by their magnitudes.
-This metric captures directional alignment in the feature space, which is often more robust to scale distortions and is considered to better reflect semantic similarity in high-dimensional embeddings (Radford et al., 2021).
+This metric captures directional alignment in the feature space, which is often more robust to scale distortions and is considered to better reflect semantic similarity in high-dimensional representations (Radford et al., 2021).
   • CKA: CKA measures similarity between two features using normalized HSIC (Hilbert-Schmidt Independence Criterion).
-Unlike MSE or cosine similarity, CKA captures structural relationships and global correlation patterns between two features, making it suitable for comparing multi-token or multi-patch feature distributions as commonly found in vision transformers.
-These metrics form diverse baselines from element-level to structure-level distortion metrics.
+It captures higher-order statistical relationships, making it particularly suitable for comparing architectural differences and global feature patterns.
 Refer to caption Refer to caption Refer to caption Refer to caption
@@
 The first, second, and third rows correspond to the Cls, Seg, and Dpt tasks, respectively.
-The first, second, third, and fourth column correspond to the HM, VTM, multi-task-trained Hyperprior, and task-specific-trained Hyperprior codecs.
+The first, second, third, and fourth columns correspond to the HM, VTM, multi-task-trained Hyperprior, and task-specific-trained Hyperprior codecs.
 5.2. Evaluation Protocol
-We define a evaluation protocol to assess how well a CFQA metric predicts the semantic distortion of compressed features. PLCC and SROCC are used in our experiments.
+We define an evaluation protocol to assess how well a CFQA metric predicts the semantic distortion of compressed features. PLCC and SROCC are used in our experiments.
   • PLCC (Linearity): PLCC quantifies how well the predicted quality scores approximate the actual semantic distortion in a linear sense.
   • SROCC (Monotonicity): SROCC assesses the consistency in the ranking between predicted and true distortions.
-Specifically, for each original feature $\mathbf{f}_{i}$, we compute these two metrics on its 10 predicted quality scores and 10 true semantic distortion.
+Specifically, for each original feature $\mathbf{f}_{i}$, we compute these two metrics on its 10 predicted quality scores and 10 true semantic distortions.
 Each metric is evaluated on all three tasks and across all codecs.
@@
 Table 2 presents the rate-accuracy performance of the handcrafted codecs.
-For all three tasks, both codecs exhibit a broad range of bitrates and corresponding accuracy levels, spanning from nearly lossless reconstruction to significant semantic degradation.
-This wide span underscores their utility in effectively simulating traditional coding distortions.
-However, the Dpt task shows a narrower variation in both bitrate and performance.
-This can be attributed to the use of multi-scale features, which include lower-level layers with higher redundancy and reduced semantic abstraction, resulting in less fluctuation in both bitrate and distortion.
+For all three tasks, both codecs exhibit a broad range of bitrates and corresponding accuracy levels, spanning from nearly lossless reconstruction to significant performance degradation.
+This wide variation highlights their effectiveness in simulating traditional coding distortions.
+However, the Dpt task exhibits less variation in both bitrate and performance.
+This difference stems from the use of multi-scale features, which include lower-level layers with higher redundancy and reduced semantic abstraction.
+These characteristics lead to smaller fluctuations in both bitrate and distortion.
 Table 3 shows the rate-accuracy performance of two learning-based codecs.
-In contrast to handcrafted codecs, the learning-based codecs exhibit more varied performance behaviors across tasks.
-Specifically, the multi-task codec outperforms the task-specific codec in Cls, whereas the task-specific codec demonstrates superior performance in Dpt.
-This suggests that incorporating fine-grained task features enhances the compression training of coarse-grained features, while fine-grained features are more effectively trained using a single distribution tailored to the task.
+Unlike the handcrafted codecs, the learning-based codecs show more diverse performance patterns across tasks.
+The multi-task Hyperprior codec achieves better performance than the task-specific Hyperprior codec for Cls, while the task-specific Hyperprior codec performs better in Dpt.
+These results indicate that incorporating fine-grained task features improves the codec optimization of coarse-grained features, whereas fine-grained features benefit more from training on a single task-specific distribution.
 These four codecs comprehensively simulate a wide range of compression distortions, making them not only highly suitable but also essential for supporting research in compressed feature quality assessment.
@@
 Additionally, since handcrafted codecs are block-based, they tend to exhibit more consistent and predictable distortion patterns.
-Among the three baseline metrics, cosine similarity consistently demonstrates a higher degree of linearity and monotonicity in relation to true semantic distortion.
-This is due to the metric’s ability to capture angular relationships between feature vectors, making it particularly effective for high-dimensional feature sets such as those produced by ViT.
+Among the three baseline metrics, cosine similarity consistently demonstrates a higher degree of linearity and monotonicity in relation to ground-truth semantic distortion.
+This is due to the metric’s ability to capture angular relationships between feature vectors, making it particularly effective for high-dimensional features such as those produced by ViTs.
 For handcrafted codecs, the three baseline metrics show better performance in capturing HM-generated distortions compared to VTM.
@@
 For most codecs, the baseline metrics show poorer fitting for Seg compared to Cls and Dpt.
-This reflects the higher semantic complexity in Seg, which is more challenging for these metrics to capture accurately.
+This reflects the higher semantic complexity in Seg, which is more challenging for these metrics to capture.
 5.4.2. Distribution Analysis
@@
 Among the three baseline metrics, cosine similarity exhibits the most concentrated PLCC distributions.
-This observation aligns with its superior average PLCC values, as reported in Table 2 and Table 3, indicating that cosine similarity provides a more stable quality estimation across varying conditions.
+This observation aligns with its superior average PLCC values, as reported in Table 4, indicating that cosine similarity provides a more stable semantic distortion measurement across varying conditions.
 Across all three metrics, Seg shows more dispersed PLCC distributions compared to Cls and Dpt.
@@
 The broader distribution further confirms that existing baseline metrics struggle to model such complex semantic degradation accurately.
-5.5. Limitations
-While the evaluated metrics demonstrate moderate performance in certain tasks and codecs, they also exhibit significant limitations.
-First, signal-based metrics like MSE lack task sensitivity and often fail to reflect meaningful semantic degradation, especially in segmentation and generation tasks.
-Second, cosine and CKA metrics, although capturing structural alignment, are sometimes unstable under nonlinear distortions introduced by learned codecs.
-Finally, none of the metrics generalize consistently across all tasks and compression strategies, suggesting that handcrafted similarity alone may not suffice for universal FQA.
-These limitations highlight the need for more adaptive, task-aware, or learned semantic quality estimators, which we consider for future work.
+5.5. Discussion
+Our evaluation reveals three key limitations of current CFQA metrics.
+First, signal-based metrics like MSE show limited task sensitivity and often correlate poorly with actual semantic degradation.
+Second, while cosine similarity and CKA capture structural relationships, they demonstrate instability when handling nonlinear distortions from learned codecs.
+Most importantly, none of these metrics achieves consistent performance across all tasks and compression methods, indicating that conventional similarity measures alone cannot provide universal CFQA solutions.
+These findings underscore the necessity for developing more adaptive, task-aware quality estimators – a direction we plan to explore in future work.
 6. Conclusion
-This paper introduces the concept of Compressed Feature Quality Assessment (CFQA), a crucial area for evaluating the semantic degradation of compressed features in systems where features are transmitted, stored, and reused.
+This paper introduces the concept of Compressed Feature Quality Assessment, a crucial research area for evaluating the semantic distortion of compressed features in systems where features are transmitted, stored, and reused.
 We present the first benchmark dataset for CFQA, which lays the groundwork for further research in this field.
-We assess the widely used metric in CFQA and provide insights for interested researchers.
-Future research will focus on developing adaptive CFQA metrics that generalize across tasks and feature coding strategies.
+We assess the widely used metrics in CFQA and provide insights for interested researchers.
+Moving forward, we will focus on developing adaptive CFQA metrics capable of generalizing across diverse tasks and coding strategies.
+Acknowledgements.
+This work was supported by the Ministry of Education of Singapore under Grant T2EP20123-0006.
+We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC.